OpenStudy (anonymous):

Express the complex number in trigonometric form. 2i

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@WhateverYouSay

imqwerty (imqwerty):

a helpful link - http://www.ditutor.com/complex_numbers/trigonometric_form.html

OpenStudy (triciaal):

i = SQRT of -1

OpenStudy (triciaal):

@imqwerty yes link helpful

OpenStudy (anonymous):

\[2 \iota=0+\iota2=r \cos \theta+\iota r \sin \theta\] \[r \cos \theta=0\] \[r \sin \theta=2\] square and add \[r=2\] \[2 \sin \theta=2,\sin \theta=1=\sin \frac{ \pi }{ 2 }=\sin \left( 2 n \pi+\frac{ \pi }{ 2 } \right)\] \[\theta=2 n \pi+\frac{ \pi }{ 2 }\] \[2 \iota=2\left\{ \cos \left( 2 n \pi+\frac{ \pi }{ 2 } \right)+\iota \sin \left( 2 n \pi+\frac{ \pi }{ 2 } \right) \right\}\] n is an integer.

OpenStudy (anonymous):

thank you all the link really helped!